OpenStudy (anonymous):

If h = 6 units, X = 9 units, Y = 13 units, and Z = 18 units, what is the volume of the triangular prism shown above?

OpenStudy (anonymous):

WILL GIVE MEDAL TO THE FIRS CORRECT ANSWER PLEASE HELP!!!

OpenStudy (anonymous):

diagram plz!

OpenStudy (anonymous):

OH WAIT

OpenStudy (anonymous):

it is given shown above with out proper source it cant be solved!

OpenStudy (anonymous):

srry the diagram is above

OpenStudy (anonymous):

ok First find the area of the triangle its 1/2 *13*6=39 then multiply with z=39*18=702

OpenStudy (anonymous):

okayy thank you

OpenStudy (anonymous):

it is the volume of the prism! that u want from the diagram! i hope iam correct!